The general term of Un of a series is defined by the relation Un+1=0.3*(1+Un). n is greater than or equal to 1 and U1 =1. Calculate U2, U3 and U4.
A second series is defined with the general term Vn, by the relation Vn=Un - 0.5 n is greater than or equal to 1.
Show that this series(Vn) is an exponential series.
Express Vn and hence Un as function of n.

Answer by stanbon(75887)

The general term of Un of a series is defined by the relation Un+1=0.3*(1+Un). n is greater than or equal to 1 and U1 =1.
Calculate
U2 = 0.3(1+1) = 0.6
U3 = 0.3(1.6) = 0.48
U4 = 0.3(1.48 = 0.444
----------------------
A second series is defined with the general term Vn, by the relation Vn=Un - 0.5 n is greater than or equal to 1.
Show that this series(Vn) is an exponential series.
V1 = U(1)-0.5
V2 = [U(1)-0.5]2 = 2U(1)-1
V3 = [2U(1)-1]3 = 6U(1)-3
I see no exponentiation.
Cheers,
Stan H.
----------
-------------------

Express Vn and hence Un as function of n.